9.55 According to the Organization for Economic Cooperation and Development (OECD), the weekly hours worked in manufacturing in the United States was 42 hours in 2017. A recent random sample of 60 manufacturing workers in the United States was selected, and it was found that they worked an average of 38.6 hours last week with a sample standard deviation of 10.4 hours.

a. Using alpha=0.05, is there enough evidence from this sample to conclude that the average number of hours of manufacturing hours worked per week is different than 42 hours?

b. Determine the precise p-value for this test using Excel
Answer #1

Given:

= 42, n = 60, = 38.6, SD(S) = 10.4, = 0.05

Hypothesis:

Ho: = 42

Ha: 42

Test statistic:

Degree of Freedom = n-1 = 60 - 1 = 59

Critical value:

Conclusion:

|Test statistic| > Critical value, i.e. 2.53 > 2.001, That is Reject Ho at 5% level of signfiicance.

b)

P-value = 0.0140 .............Using t table

P-value < , i.e 0.0140 < 0.05, That is Reject Ho at 5% level of signficance.

Therefore, there enough evidence from this sample to conclude that the average number of hours of manufacturing hours worked per week is different than 42 hours